OpenStudy (anonymous):

how you the parabola graph of given function with 4 points (2,0),(3.2)

OpenStudy (anonymous):

since y=ax2+bx+c, for the first point: 0=4a+2b+c and for the second point: 2=9a+3b+c. the equations are less than the variables. the system is consistent/have at least 1 soluiton.